OpenStudy (anonymous):

When a positive integer p is divided by 7, the remainder is 2. Which of the following expression will yield a remainder of 4 when divided by 7?

OpenStudy (anonymous):

1.p+2 2.P+3 3.P+4 4.P+5 5.P+6

ganeshie8 (ganeshie8):

any ideas ?

ganeshie8 (ganeshie8):

maybe use your knowledge about long division and make a best guess...

OpenStudy (anonymous):

... I would guess its P+2 im not sure...

jimthompson5910 (jim_thompson5910):

`a positive integer p is divided by 7, the remainder is 2` means we have this equation p/7 = q + 2/7 where q is some integer (q for quotient)

jimthompson5910 (jim_thompson5910):

multiply both sides by 7 and distribute to get p = 7q + 2

jimthompson5910 (jim_thompson5910):

the goal from here is to somehow get that '2' to turn into '4' we do this by adding 2 to each side, so, p = 7q + 2 p+2 = 7q + 2+2 p+2 = 7q + 4 that turns back into (p+2)/7 = q + 4/7 so you're right, the answer is indeed p+2